OpenStudy (anonymous):

for n>= 1 use mathematical induction

OpenStudy (anonymous):

\[15/2^{4n} -1\]

OpenStudy (anonymous):

i got most of it im just stuck..

OpenStudy (anonymous):

this is what i have

OpenStudy (anonymous):

can you state the complete question...

OpenStudy (anonymous):

yes give me a second

ganeshie8 (ganeshie8):

\(15 | (2^{4n} -1) \)

OpenStudy (anonymous):

\[15|2^{4n}-1=15p\] \[P(1): 15|2^{4(1)}-1\] ==> \[15|2^4-1\] =>\[15|15 true \] P(k)\[15|2^{4k}-1<=>2^{4k}-1=15p\] P(k+1)\[15|2^{4k+4}-1<=>2^{4k+4}-1=15m\] \[2^4*2^k-1=15m\]

OpenStudy (anonymous):

so then what it becomes? \[16(2^{4k}-1=15m\]

ganeshie8 (ganeshie8):

Looks good !

OpenStudy (anonymous):

ok but im stuck right here...

ganeshie8 (ganeshie8):

you multiplied 2^4 both sides, eh ?

OpenStudy (anonymous):

now we need to prove on that 15 divides 2^(4*(k+1) - 1 using 2^4k -1 =15p

OpenStudy (anonymous):

ok...

ganeshie8 (ganeshie8):

P(k) : \(2^{4k}-1=15p \) multiply \(2^4\) both sides : \(2^4\left(2^{4k}-1\right)=15q \) \( 2^{4k+4}-16=15q \) \(2^{4(k+1)}-1-15=15q \)

OpenStudy (anonymous):

2^(4*(k+1) - 1 =2^4 *2^4k - 1 =2^4(15p+1) -1

OpenStudy (zarkon):

\[2^{4(k+1)}-1=2^{4k}2^4-1=2^{4k}(2^4-1+1)-1\] \[=2^{4k}(2^4-1)+2^{4k}-1\]

OpenStudy (anonymous):

=15*16p +16 -1 = 15*16p +15=15(16p+1)

OpenStudy (anonymous):

p is integer hence (16p+1) is also an integer

OpenStudy (anonymous):

hence further...

Miracrown (miracrown):

2^[4*(k+1)] = 2^(4k+4) = 2^4 * 2^(4k) So now we have to show that 2^4 * 2^(4k) - 1 = 15m 2^4 * 2^(4k) - 1 = 2^4 * 2^(4k) - 16 + 15 = 16 * [2^(4k) - 1] + 15 = 16 * 15p + 15 = 16 * 15p + 15 = 15*m

Miracrown (miracrown):

I've showed that it lead to a multiple of 15 Start with 2^4 * 2^(4k) - 1 next rewrite the -1 as -16 + 15 then 2^4 * 2^(4k) - 16 + 15 2^4 = 16 So we get 16 * 2^(4k) - 16 + 15 Next we factor 16 from the first two terms 16 [2^(4k) - 1] + 15 Now the thing in square brackets is our assumption [2^(4k) - 1] = 15p So substituting that we get 16*15p + 15 Next we factor out the 15 15*[16p - 1] So since 16p-1 is also an integer, I have shown through algebra that 2^[4(k+1)] - 1 = 15*[16p -1] = 15 * m therefore 15 divides 2^[4(k+1)] - 1 where m is the integer equal to 16p-1

OpenStudy (anonymous):

thanks for the detail thats what i needed! lol

Miracrown (miracrown):

ha ha, no worries. :-]